# An Upper Bound on the Error Probability of RPA Decoding of Reed-Muller Codes Over the BSC

### v1

@@ VI Conclusion and Future Work
In this paper, we provided a formal, theoretical analysis of the performance of the Recursive Projection-Aggregation (RPA) decoder for Reed-Muller (RM) codes over a binary symmetric channel (BSC). Via our analysis, we obtained that the RPA decoder recovers the correct input codeword with overwhelming probability, for RM codes of order logarithmic in $m$, for large enough blocklengths $N=2^{m}$. Key components of our analysis were our estimates of the error probabilities of the Fast Hadamard Transform-based decoder and the aggregation step in the RPA decoder.
@@
An interesting line of future work would be to extend the analysis in this paper to situations where more than one iteration of the RPA decoder is run, before convergence. We believe that such an attempt will allow one to obtain vanishing error probabilities for higher RM orders than guaranteed in this paper. Another important extension of our analysis will be to the setting of the transmission of RM codewords over general binary-input, memoryless, symmetric (BIMS) channels (or even over general discrete memoryless channels or DMCs), of which the BSC is a special case.
@@ Acknowledgement
This work was carried out when the second author was visiting the Indian Institute of Science, Bengaluru. The authors thank Prof. Navin Kashyap and Prof. Henry Pfister for helpful comments and suggestions.

### v2

@@ VII Conclusion and Future Work
In this paper, we provided a formal, theoretical analysis of the performance of the Recursive Projection-Aggregation (RPA) decoder for Reed-Muller (RM) codes over a binary symmetric channel (BSC). Via our analysis, we obtained that the RPA decoder recovers the correct input codeword with overwhelming probability, for RM codes of order logarithmic in $m$, when the blocklength $N=2^{m}$ is sufficiently large. Key components of our analysis were our estimates of the error probabilities of the Fast Hadamard Transform-based decoder and the aggregation step in the RPA decoder.
@@
An interesting line of future work would be to extend the analysis in this paper to situations where more than one iteration of the RPA decoder is run, before convergence. We expect that such an attempt will allow one to obtain vanishing error probabilities for higher RM orders than guaranteed in this paper. Another important extension of our analysis will be to the setting of the transmission of RM codewords over general binary-input, memoryless, symmetric (BIMS) channels (or even over general discrete memoryless channels or DMCs), of which the BSC is a special case.
@@ Acknowledgement
This work was carried out in part when the second author was visiting the Simons Institute for the Theory of Computing, Berkeley, during Jan-May 2024 and the Indian Institute of Science, Bengaluru, during Aug-Dec 2024. The authors thank Ilya Dumer, Navin Kashyap, and Henry Pfister for helpful comments and pointers to the literature.